x^4 = 1.02, how do i solve for x? and how do i check it if it is correct?
Answer by rothauserc(4718)

we take the fourth root of both sides of the =
x = (1.02)^(1/4)
x = 1.005 or -1.005
check the answer
(1.005)^4 = 1.02
(-1.005)^4 = 1.02